OpenStudy (anonymous):

it seems that my math problems are confusing me, i thank every1 tats helping me with slope intercept but whats an easier way to solve slope problems? what if i have problems like x - y + 1= 0? or 2x + 4y = 5? can i get an example using these two problems

OpenStudy (anonymous):

these are all linear equations, so the simple way to find slope is to put y in one side and everything in the other side. x-y+1=0 -> y=x+1, so slope is 1 (remember 'y=mx + a' where m is slope and a is a number) similarly, 2x + 4y = 5 -> 4y= -2x + 5 -> y=(-2/4)x + 5/4 so m=-2/4=-1/2

OpenStudy (anonymous):

ohh ok, i see now , so were just using y=mx+b and like u said 'a" would be the number. or in this case 'b' would be the number

OpenStudy (anonymous):

are we just plugging the equation in, were not changing anything ?

OpenStudy (anonymous):

yeah, a, b can be anything, not important at all when you find the slope. and yes, just change the equation a bit so that it is in the form of y=mx + whateveritis. for example if you have 99y, then divide every terms in the equation for 99, so we can have y stand a lone.

OpenStudy (anonymous):

oh oki see, well thanks alot =)

OpenStudy (anonymous):

no prob :D